OpenStudy (anonymous):

what expression is equivalent to m/n to the 7th power

OpenStudy (lgbasallote):

this is \[\huge (\frac mn)^7\]right? here's a hint... \[\huge (\frac xy)^2 \implies \frac{x^2}{y^2}\] does that help?

OpenStudy (anonymous):

yes thank you!

OpenStudy (lgbasallote):

welcome ^_^